Identify the vertex, focus, axis of symmetry, directrix, direction of opening, length of the latus rectum, and the x- and y-intercepts of: 2y ^ 2 + x + 20y + 51 = 0


1
Expert's answer
2021-11-24T17:59:58-0500
"2y ^ 2 + x + 20y + 51 = 0"

"x=-2(y ^ 2 +10y+25) -1"

The equation of the parabola in vertex form:


"x=-2(y+5)^2-1"

Vertex: "(-1, -5)"


"(y+5)^2=4(-\dfrac{1}{8})(x+1)"

Focus: "(-1-\dfrac{1}{8}, -5)=>"Focus: "(-\dfrac{9}{8}, -5)."

Axis of symmetry: "y=-5."

Directrix: "x=-1-(-\dfrac{1}{8})=>" Directrix: "x=-\dfrac{7}{8}."

The horizontal parabola that opens to the left.

The length of the latus rectum of the parabola is "\dfrac{1}{2}."

x-intercept: "(\u221251,0)"

No y-intercepts.